Please help me solve these problem.
Ignoring air resistance, the distance d in feet that an object falls in t seconds is given by the formula d=1/2gt^2, where g is approximately 32 ft.per second sq. this formula can also be express as d=16t^2.
a. A plant falls from a window ledge 144 feet high. How long does it takes to hit the street below?
b. A piece of concrete falls from a bridge 400 feet high,how long does it takes to hit the water below?



Answer by stanbon(57967)

Ignoring air resistance, the distance d in feet that an object falls in t seconds is given by the formula d = (1/2)gt^2, where g is approximately 32 ft.per second sq. this formula can also be express as d=16t^2.
a. A plant falls from a window ledge 144 feet high. How long does it takes to hit the street below?
Solve: 144 = 16*t^2
t = sqrt(144/16) = 12/4 = 3 seconds
------------------------------------------
b. A piece of concrete falls from a bridge 400 feet high,how long does it takes to hit the water below?
Solve: 400 = 16t^2
t^2 = (400/16) = 20/4) = 5 seconds
====================================
Cheers,
Stan H.
===================